OpenStudy (anonymous):

Solve: 5(x + 8) = 7(x − 2) 17 −17 27 −27

OpenStudy (anonymous):

27.

OpenStudy (anonymous):

5(27 + 8) = 175 7(27 - 2) = 175

OpenStudy (anonymous):

5(x + 8) = 7(x − 2) 5x + 40 = 7x − 14 40+14= 7x-5x 54=2x x=54/2 x=27